OpenStudy (anonymous):

Factor 9x 6 – 16y 6 completely.

OpenStudy (anonymous):

this doesn't make sense

OpenStudy (anonymous):

9x^6 - 16y^6

OpenStudy (anonymous):

ah there you go

OpenStudy (anonymous):

(3x^6 -4y^3)(3x^6 +4y^3)

OpenStudy (anonymous):

\[(3x^{6}-4y^{3})(3x ^{6}+4y ^{3})\]

OpenStudy (anonymous):

9x^6 – 16y^6 =

OpenStudy (anonymous):

my equation is your answer

OpenStudy (anonymous):

ohhhh, Thanks

OpenStudy (anonymous):

multiply it out and you'll see

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you help me with this? Factor x^4 + x y^3 + x^3 + y^3 completely. Show your work. Hint: First factor out a monomial.

hero (hero):

The formula for difference of squares is \(a^2 - b^2 = (a + b)(a - b)\) In this case \(9x^6 - 16y^6 = (3x^3 - 4y^3)(3x^3 + 4y^3)\) Check: \((3x^3 - 4y^3)(3x^3 + 4y^3) = (3x^3(3x^3 + 4y^3) - 4y^3(3x^3 + 4y^3)\) \(=(9x^6 + 12x^3y^3 - 12x^3y^3 - 16y^6)\) \(=9x^6 - 16y^6\)